OpenStudy (anonymous):

What is the expression in factored form? x^2+ 13x + 42 help please?

hartnn (hartnn):

can you tell me 2 numbers whose sum is 13 and product = 42 ?

OpenStudy (anonymous):

7 and 6

OpenStudy (anonymous):

so the answer would be (x+7)(x+6)

hartnn (hartnn):

let lexi try ?

OpenStudy (anonymous):

6 and 7 6+7 = 13 6*7 = 42 :)

hartnn (hartnn):

yes! do you know how to factor now ?

OpenStudy (anonymous):

not exactally..:/

hartnn (hartnn):

ok, let me tell u so we split here 13 x = 6x+7x we have (x^2+6x)+(7x +42) ok ?

OpenStudy (anonymous):

ok

hartnn (hartnn):

what can you factor out from 1st 2 terms ? x^2+6x =... ?

OpenStudy (anonymous):

7x + 42?

hartnn (hartnn):

lets consider 1st 2 terms initially x^2+6x = x*x+6*x what do you see common in these 2 ?

OpenStudy (anonymous):

x*x is the same as x^2? and 6*x is the same as 6x

hartnn (hartnn):

no, i meant 'x' is common from x*x+ 6x so, x^2+6x = x (x+6) got this ?

OpenStudy (anonymous):

ooh ok:)

hartnn (hartnn):

so, what do you see common from last 2 terms 7x+42 ?

OpenStudy (anonymous):

well.. if u divide 42 by 7 u get 6.. but thats about it

hartnn (hartnn):

see, 7x +42 = 7*x +7*6 so, 7 is common, so i take that out 7x+42 = 7(x+6) makes sense ?

OpenStudy (anonymous):

oooh yea ok i see what u did:)

OpenStudy (anonymous):

thank you!!:)

hartnn (hartnn):

good :) so, we have x(x+6) +7 (x+6) now what is common in these 2 terms ?

OpenStudy (anonymous):

is that all we do?

hartnn (hartnn):

last step is remaining.... whats common term in x(x+6) +7 (x+6) ?

OpenStudy (anonymous):

the (x+6)'s?

hartnn (hartnn):

yes! so factor that out! what remains is x and +7 so, x(x+6) +7 (x+6) = (x+6)(x+7) thats it!

OpenStudy (anonymous):

thank you sooo much!! so the answers would be x+6 and x+7?:)

hartnn (hartnn):

the answer = (x+6)(x+7) no 'and' in between welcome ^_^